OpenStudy (anonymous):

A car traveled 75% of the way from town A to town B at an average speed of 50 kmph. The car travels at an average speed of S kmph for the remaining part of the trip. The average speed for the entire trip was 40 kmph. What is S?

OpenStudy (anonymous):

x / [(0.75x)/50 + (0.25x)/s] = 40 x actually cancels out. s = 25 km/hr

OpenStudy (anonymous):

\[2\times50\times S \div 50+ S =40\] remaining 25 % will be covered in 33.33 kmph